- Press Release
- August 15, 2022
NOAA GOES SATOPS Morning Report 24 August 2000
SATOPS MORNING REPORT August 24, 2000
No change in the status of the POES spacecraft.
NOAA-14 rev 29112F – G1A noisy from spacecraft with 92.6% of the data
received. Recorder management prevented replay from the spacecraft on subsequent
NOAA-10 rev 72443 and 72444/ F – Executed a series of manual commands
to offset the solar array from -65 to -72 degrees. All sub-system
telemetry appears nominal, including ADACS and power.
POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
An OPS48 Category III data loss on rev 17476 was declared by AFWA.
The data would not process on AFWA equipment (reasons unknown)
OPS-46 deactivation activities continue. The solar array lead
angle adjustment testing continues (yesterday and today) to evaluate effects
on the Battery 2 Current Sensor (BCS).
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
At 1853z on Wednesday, August 23, a GOES-10 frame break was observed
and attributed to high BER (unknown cause).
At 0512z on Thursday, August 24, a GOES-10 frame break was observed
and attributed to loss of signal synchronization.
Normal pre-eclipse battery reconditioning activities continue on GOES-8,
-9 and -10.
GOES operations were nominal over the past 24 hours.
Questions or comments to firstname.lastname@example.org